Grade 12th passAlgebra

(a+b+c)^2

Sarnali sarkar
9 Years agoGrade 12th pass

1 Answer

Tanishq Chourishi
9 Years ago

Now we will learn to expand the square of a trinomial (a + b + c).

Let (b + c) = x      

 

Then (a + b + c)2 = (a + x)2 = a2 + 2ax + x2

                         = a2 + 2a (b + c) + (b + c)2 

                         = a2 + 2ab + 2ac + (b2 + c2 + 2bc) 

                         = a2 + b2 + c2 + 2ab + 2bc + 2ca

Therefore, (a + b + c)2 = a2 + b2 + c2 + 2ab + 2bc + 2ca